OpenStudy (anonymous):

solve the trigonometric equation. tan (pi/2-x) + tan (-x)= 0

OpenStudy (anonymous):

is \[\tan (\frac{ \pi }{ 2}-x)+\tan(x) or \tan(\frac{ \pi }{ 2-x })+\tan(x)?\]?

OpenStudy (anonymous):

if is the firts, the answer is X=pi/4 but if is the second the answer is a little complicated.

OpenStudy (anonymous):

it is the first and I know the answer but i want to know how to do it

OpenStudy (anonymous):

well first apply tan^-1(argument) or arctan that is the reverse of tangent, then \[\tan^-1(\tan(\frac{ \pi }{ 2 }-x))=\tan^-1(\tan(x))\] \[\frac{ \pi }{ 2 }-x=x \] so u can clear the x\[\frac{ \pi }{ 2 }=2x \] then x=pi/4